# Solving Optimal Power Flow using a Variational Quantum Approach

### v1

@@ 3 Primal-dual iterations for the OPF
Although (QCQP) is nonconvex, it can be relaxed to an SDP and solved by standard interior-point solvers. For many practical settings, this convex relaxation is successful in the sense that a globally optimal $\mathbf{v}$ can be readily recovered from the SDP solution. Nonetheless, the memory and time requirements imposed by SDP solvers may be challenging to meet as $N$ increases. This motivates us to explore a VQA for the OPF.
@@
To motivate a PQC-based solver for large-scale OPFs, let us discuss how (QCQP) could be solved on a classical computer. Among different approaches, we present the method of primal-dual (PD) decomposition as a representative example of a first-order optimization algorithm. The PD method relies on Lagrangian duality [3]. Each constraint indexed by $m$ in (QCQP) is associated with a non-negative dual variable $\lambda_{m}$. Let a vector $\boldsymbol{\lambda}\in\mathbb{R}^{M}_{+}$ collect all dual variables. The related Lagrangian function is a function of both $\mathbf{v}$ and $\boldsymbol{\lambda}$:
$\mathcal{L}(\mathbf{v};\boldsymbol{\lambda})=\mathbf{v}^{\dagger}\mathbf{M}_{0}\mathbf{v}+\sum_{m=1}^{M}\lambda_{m}(\mathbf{v}^{\dagger}\mathbf{M}_{m}\mathbf{v}-b_{m}).$ (1)
@@
$D^{*}\coloneqq\max_{\boldsymbol{\lambda}\succeq\mathbf{0}}\min_{\mathbf{v}\in\mathbb{C}^{N}}\leavevmode\nobreak\ \mathcal{L}(\mathbf{v};\boldsymbol{\lambda}).$ (2)
A primal/dual solution $(\mathbf{v}^{*},\boldsymbol{\lambda}^{*})\in\mathbb{C}^{N}\times\mathbb{R}^{M}_{+}$ of (2) satisfies
$\mathcal{L}(\mathbf{v}^{*};\boldsymbol{\lambda})\leq\mathcal{L}(\mathbf{v}^{*};\boldsymbol{\lambda}^{*})\leq\mathcal{L}(\mathbf{v};\boldsymbol{\lambda}^{*})\quad\leavevmode\nobreak\$
for all $(\mathbf{v},\boldsymbol{\lambda})\in\mathbb{C}^{N}\times\mathbb{R}^{M}_{+}$. Such a pair is called a saddle point of the Lagrangian function. In essence, $\mathbf{v}^{*}$ is a minimizer of $\mathcal{L}(\mathbf{v};\boldsymbol{\lambda}^{*})$ and $\boldsymbol{\lambda}^{*}$ is a maximizer of $\mathcal{L}(\mathbf{v}^{*};\boldsymbol{\lambda})$. For problems featuring strong duality (in which case $P^{*}=D^{*}$), a point is an optimal primal/dual point if and only if it is a saddle point of the Lagrangian function [3, p. 239]. Furthermore, at a primal/dual solution $(\mathbf{v}^{*};\boldsymbol{\lambda}^{*})$, the value of the Lagrangian satisfies $\mathcal{L}(\mathbf{v}^{*};\boldsymbol{\lambda}^{*})=P^{*}=D^{*}$. Finding an exact saddle point of the Lagrangian imposes stringent conditions on $\mathcal{L}(\mathbf{v};\boldsymbol{\lambda})$. However, approximate saddle points can be reached under specific settings using gradient-based methods, such as the PD method [37].
@@
$\mathbf{v}^{t+1}$ $\coloneqq\mathbf{v}^{t}-\mu^{t}_{v}\nabla_{\mathbf{v}}\mathcal{L}(\mathbf{v}^{t};\boldsymbol{\lambda}^{t}),$ (3a)
@@
where $\mu_{v}^{t}$ and $\mu_{\lambda}^{t}$ are positive step sizes, and $x]_{+}\coloneqq\max\{x,0\}$. The update in (3a) is a gradient descent step on $\mathcal{L}$ in terms of the primal variable. The update in (3b) is a projected gradient ascent step on $\mathcal{L}$ in terms of the dual variable.
One may attempt to solve (QCQP) using the previous PD iterations. In this case, computing the Lagrangian gradients incurs complexity $\mathcal{O}(N)$. To see this, note that the partial derivative of $\mathcal{L}$ with respect to $\lambda_{m}$ is
$\frac{\partial\mathcal{L}}{\partial\lambda_{m}}=\mathbf{v}^{\dagger}\mathbf{M}_{m}\mathbf{v}-b_{m}.$
This partial derivative can be computed in $\mathcal{O}(1)$ steps, simply because $\mathbf{M}_{m}$ has $\mathcal{O}(1)$ nonzero entries. Therefore, computing the entire gradient $\nabla_{\boldsymbol{\lambda}}\mathcal{L}$ entails complexity $\mathcal{O}(M)=\mathcal{O}(N)$ given that $M$ is proportional to $N$. Similarly, computing the gradient
$\nabla_{\mathbf{v}}\mathcal{L}=2\left(\mathbf{M}_{0}+\sum_{m=1}^{M}\lambda_{m}\mathbf{M}_{m}\right)\mathbf{v}$
incurs $\mathcal{O}(N)$ operations. Contemporary power systems consist of more than $N=30,000$ nodes. Nonetheless, it is anticipated that the size of the OPF could increase dramatically due to the pressing need to schedule renewables at finer spatiotemporal scales. In this context, can a VQA attain reduced computational complexity? The next section proposes a variational model for the OPF.
4 A doubly variational OPF model
Variational quantum computing is an algorithmic tool for solving high-dimensional problems in a parameterized form using a parameterized quantum circuit (PQC). The well-known variational quantum eigensolver (VQE) solves a specific form of (QCQP[13]:
$\min_{\mathbf{v}\in\mathbb{C}^{N}}\leavevmode\nobreak\$$\leavevmode\nobreak\ \mathbf{v}^{\dagger}\mathbf{M}_{0}\mathbf{v}$ (4)
s.to $\leavevmode\nobreak\ \mathbf{v}^{\dagger}\mathbf{v}=1.$
@@ 4 A doubly variational OPF model
We propose solving the OPF using a similar workflow. Because the primal variable of the OPF does not have a unit Euclidean norm, we parameterize it as [38]:
@@
$P_{\theta}^{*}\coloneqq\min_{\boldsymbol{\theta},\alpha>0}\leavevmode\nobreak\$$\leavevmode\nobreak\ \alpha^{2}F_{0}(\boldsymbol{\theta})$ (QCQPθ)
$\mathrm{s.to}\leavevmode\nobreak\$$\leavevmode\nobreak\ \alpha^{2}F_{m}(\boldsymbol{\theta})\leq b_{m},\quad m=1:M$
@@
As in VQE, the original (QCQP) over $\mathbf{v}$ has been replaced by the variational problem (QCQPθ) over the parameter vector $\boldsymbol{\theta}$. Attempting to solve (QCQPθ) using Lagrangian duality entails solving the dual problem:
$D_{\theta}^{*}\coloneqq\max_{\boldsymbol{\lambda}\geq\mathbf{0}}\leavevmode\nobreak\ \min_{\boldsymbol{\theta},\alpha>0}\mathcal{L}(\boldsymbol{\theta},\alpha;\boldsymbol{\lambda})$
@@
$\lambda_{m}(\boldsymbol{\phi})=\beta^{2}\cdot|\xi_{m}(\boldsymbol{\phi})|^{2},\quad\text{for}\leavevmode\nobreak\ m=1:M$ (8)
@@
The Lagrangian function related to this doubly parameterized OPF can be expanded into three terms:
@@
5 Solving the doubly variational OPF
@@
Lemma 1 asserts that $F(\boldsymbol{\theta},\boldsymbol{\phi})$ is an expectation applied to the composite state of the two PQCs. This is important as it bypasses the need to measure separately each one of the $M$ constraint expectations $F_{m}(\boldsymbol{\theta})$ of the OPF. This powerful feature was first identified in [18] for solving doubly variational SDPs on a quantum computer. In this work, we adopt this idea to the OPF setting. We also propose a method for efficiently measuring the joint observable for the OPF associated with sparse graphs in general. We uniquely leverage the graph structure of a power system to efficiently measure the three observables using the two PQCs. We defer this discussion to Section 6. If the observables can be measured efficiently, their gradients can be measured efficiently as well via the so-called parameter-shift rule (PSR). We review this rule and adapt it to the PD iterations in (13).
Let us focus on the gradient $\nabla_{\boldsymbol{\theta}}F_{0}(\boldsymbol{\theta}^{t})$. Suppose that the unitary matrix modeling the primal PQC takes the form
@@
The gradient of the expectation value in (15) with respect to $\boldsymbol{\phi}$ can be computed similarly by running PQCd by shifting the values of $\boldsymbol{\phi}^{t}$ for $2Q$ times or $2QS$ measurement samples. The gradient $\nabla_{\boldsymbol{\theta}}F(\boldsymbol{\theta}^{t},\boldsymbol{\phi}^{t})$ can be computed by running both PQCs while only shifting the entries of $\boldsymbol{\theta}^{t}$; and vice versa for computing the gradient $\nabla_{\boldsymbol{\phi}}F(\boldsymbol{\theta}^{t},\boldsymbol{\phi}^{t})$. It should be clear by now that, due to the PSR, evaluating gradients of expectations is equivalent to measuring expectations. Section 6 discusses how to measure the observables involved in the OPF.
Although the plain PD method has been extensively studied for convex/concave and nonconvex/concave saddle-point problems [41, 42, 43, 44], its convergence has not been fully understood in the nonconvex/nonconcave setting, like the one in (12). Recently, a variation of the PD method, termed the extragradient (EG) method [45], has been shown to converge under certain conditions for nonconvex/nonconcave saddle-point problems [46]. Compared to the plain PD method, each EG iteration doubles the number of gradient evaluations of the Lagrangian function. More specifically, let vector $\mathbf{z}^{t}\coloneqq[\boldsymbol{\theta}^{t};\alpha^{t};\boldsymbol{\phi}^{t};\beta^{t}]$ collect the primal and dual variables at iterate $t$, and define the vector:
@@
for a step size $\mu_{z}^{t}>0$. In essence, the EG method first computes an intermediate point $\bar{\mathbf{z}}^{t}$ based on the plain PD method, and then updates $\mathbf{z}^{t}$ to $\mathbf{z}^{t+1}$ upon evaluating the gradient operator at the intermediate point. Despite doubling the number of gradient evaluations, EG iterates feature favorable convergence properties as established analytically in Section 7 and corroborated numerically in Section 8.
Before concluding this section, it is worth noting that, as with other VQAs, training the primal-dual PQC pair could be impeded by the so-called barren plateaus issue, according to which gradients vanish exponentially fast in terms of the number of qubits and depth of PQCs [47, 48]. This phenomenon can occur when designing VQAs for large-scale problems and/or using overly expressive PQCs. A related concern is that the PQC architectures that are both sufficiently expressive and free from barren plateaus are often efficiently simulated by classical algorithms [49]. Interestingly, recent results show that the so-called dynamic quantum phase gates (PQCs) [50], which incorporate early measurements followed by measurement-conditional unitaries, and quantum recurrent embedding neural networks [51] can provably avoid barren plateaus while retaining high expressivity. Moreover, dynamic PQCs inherit the worst-case classical hardness of the universal unitary circuits, precluding the possibility of classical simulatability. Investigating the performance of the doubly variational OPF on dynamic PQCs or quantum recurrent embedding neural networks for large-scale networks is an interesting future direction.
6 Measuring OPF observables
This section reviews the extended Bell measurement (XBM) method proposed in [4] and explains how it can be adapted to efficiently measure OPF observables.
6.1 Extended Bell measurement (XBM) method
@@
$\braket{i|\mathbf{M}_{m}^{c}|j}=\begin{cases}\braket{i|\mathbf{M}_{m}|j},&\text{if}\leavevmode\nobreak\ i\oplus j=c\\ 0,&\text{otherwise}.\end{cases}$
@@
We next describe how to effect an OPF measurement matrix $\mathbf{M}$ that uses a few colors.
6.2 Node permutation for qubit-efficient measurements

### v2

@@ II Primal-dual iterations for QCQP
Although (QCQP) is generally nonconvex, it can be relaxed to an SDP, which can be solved by standard interior-point solvers. In some application domains, this convex relaxation is successful in the sense that the optimal SDP variable is rank-one and a globally optimal $\mathbf{x}$ can be readily recovered from the SDP solution. Nonetheless, the memory and time requirements imposed by SDP solvers may be challenging to meet as $N$ increases. This motivates us to explore a VQA for QCQP/SDPs.
@@
Toward developing a PQC-based solver for large-scale QCQP/SDPs, let us discuss how (QCQP) could be solved on a classical computer. Among the different approaches, we present the primal-dual (PD) decomposition as a representative example of a first-order optimization algorithm. The PD method relies on Lagrangian duality [8]. Each constraint indexed by $m$ in (QCQP) is associated with a non-negative dual variable $\lambda_{m}$. Let vector $\boldsymbol{\lambda}\in\mathbb{R}^{M}_{+}$ collect all dual variables. The related Lagrangian function is a function of both $\mathbf{x}$ and $\boldsymbol{\lambda}$:
$\mathcal{L}(\mathbf{x};\boldsymbol{\lambda})=\mathbf{x}^{\dagger}\mathbf{M}_{0}\mathbf{x}+\sum_{m=1}^{M}\lambda_{m}(\mathbf{x}^{\dagger}\mathbf{M}_{m}\mathbf{x}-b_{m}).$ (1)
@@
$D^{*}\coloneqq\max_{\boldsymbol{\lambda}\succeq\mathbf{0}}\min_{\mathbf{x}\in\mathbb{C}^{N}}\penalty 10000\ \mathcal{L}(\mathbf{x};\boldsymbol{\lambda}).$ (2)
A primal/dual solution $(\mathbf{x}^{*},\boldsymbol{\lambda}^{*})\in\mathbb{C}^{N}\times\mathbb{R}^{M}_{+}$ of (2) satisfies
$\mathcal{L}(\mathbf{x}^{*};\boldsymbol{\lambda})\leq\mathcal{L}(\mathbf{x}^{*};\boldsymbol{\lambda}^{*})\leq\mathcal{L}(\mathbf{x};\boldsymbol{\lambda}^{*})\quad\penalty 10000\$
for all $(\mathbf{x},\boldsymbol{\lambda})\in\mathbb{C}^{N}\times\mathbb{R}^{M}_{+}$. Such a pair is called a saddle point of the Lagrangian function. In essence, $\mathbf{x}^{*}$ is a minimizer of $\mathcal{L}(\mathbf{x};\boldsymbol{\lambda}^{*})$ and $\boldsymbol{\lambda}^{*}$ is a maximizer of $\mathcal{L}(\mathbf{x}^{*};\boldsymbol{\lambda})$. For problems featuring strong duality (in which case $P^{*}=D^{*}$), a point is an optimal primal/dual point if and only if it is a saddle point of the Lagrangian function [8, p. 239]. Furthermore, at a primal/dual solution $(\mathbf{x}^{*};\boldsymbol{\lambda}^{*})$, the value of the Lagrangian satisfies $\mathcal{L}(\mathbf{x}^{*};\boldsymbol{\lambda}^{*})=P^{*}=D^{*}$. Finding an exact saddle point of the Lagrangian imposes stringent conditions on $\mathcal{L}(\mathbf{x};\boldsymbol{\lambda})$. However, approximate saddle points can be reached under specific settings using gradient-based methods, such as the PD method [48].
@@
$\mathbf{x}^{t+1}$ $\coloneqq\mathbf{x}^{t}-\mu^{t}_{x}\nabla_{\mathbf{x}}\mathcal{L}(\mathbf{x}^{t};\boldsymbol{\lambda}^{t}),$ (3a)
@@
where $\mu_{x}^{t}$ and $\mu_{\lambda}^{t}$ are positive step sizes, and $\lambda]_{+}\coloneqq\max\{\lambda,0\}$. The update in (3a) is a gradient descent step on $\mathcal{L}$ in terms of the primal variable. The update in (3b) is a projected gradient ascent step on $\mathcal{L}$ in terms of the dual variable.
One may attempt to solve (QCQP) using the PD iterations in (3). In this case, computing the Lagrangian gradients incurs complexity $\mathcal{O}(sN^{2})$, assuming each row of $\mathbf{M}_{m}$ has at most $s$ nonzero entries. To see this, note that the partial derivative of $\mathcal{L}$ with respect to $\lambda_{m}$ is
$\frac{\partial\mathcal{L}}{\partial\lambda_{m}}=\mathbf{x}^{\dagger}\mathbf{M}_{m}\mathbf{x}-b_{m},$
which can be computed in $\mathcal{O}(sN)$ operations. Therefore, computing the entire gradient $\nabla_{\boldsymbol{\lambda}}\mathcal{L}$ entails complexity $\mathcal{O}(sNM)=\mathcal{O}(sN^{2})$ given that $M$ is proportional to $N$. Similarly, computing the gradient
$\nabla_{\mathbf{x}}\mathcal{L}=2\left(\mathbf{M}_{0}+\sum_{m=1}^{M}\lambda_{m}\mathbf{M}_{m}\right)\mathbf{x}$
incurs $\mathcal{O}(sN^{2})$ operations. For large-scale problems, such as solving AC OPF for contemporary power systems with more than $N=50,000$ nodes, this computational cost quickly becomes prohibitive. In this context, can a VQA attain reduced computational complexity? The next section proposes a variational model for QCQP/SDPs.
III A doubly variational QCQP/SDP
Variational quantum computing is an algorithmic tool for solving high-dimensional problems in a parameterized form using a PQC. The well-known VQE solves a specific form of (QCQP) [52]:
$\min_{\mathbf{x}\in\mathbb{C}^{N}}\penalty 10000\$$\penalty 10000\ \mathbf{x}^{\dagger}\mathbf{M}_{0}\mathbf{x}$ (4)
s.to $\penalty 10000\ \mathbf{x}^{\dagger}\mathbf{x}=1.$
@@ III A doubly variational QCQP/SDP
We propose solving the (QCQP) in a similar manner. Because the primal variable of the (QCQP) does not have a unit Euclidean norm, we parameterize it as [57]:
@@
$P_{\theta}^{*}\coloneqq\min_{\boldsymbol{\theta},\alpha>0}\penalty 10000\$$\penalty 10000\ \alpha^{2}F_{0}(\boldsymbol{\theta})$ (QCQPθ)
$\mathrm{s.to}\penalty 10000\$$\penalty 10000\ \alpha^{2}F_{m}(\boldsymbol{\theta})\leq b_{m},\quad m=1:M$
@@
These expectations can be alternatively expressed via the density operator $\boldsymbol{\rho}(\boldsymbol{\theta})$ associated with the PQCp state as
$F_{m}(\boldsymbol{\theta})=\operatorname{Tr}(\mathbf{M}_{m}\boldsymbol{\rho}(\boldsymbol{\theta})),\quad m=0:M.$ (7)
[…]
As in VQE, the original (QCQP) over $\mathbf{x}$ has been replaced by the variational problem (QCQPθ) over the parameter vector $\boldsymbol{\theta}$. Attempting to solve (QCQPθ) using Lagrangian duality entails solving the dual problem:
$D_{\theta}^{*}\coloneqq\max_{\boldsymbol{\lambda}\geq\mathbf{0}}\penalty 10000\ \min_{\boldsymbol{\theta},\alpha>0}\mathcal{L}(\boldsymbol{\theta},\alpha;\boldsymbol{\lambda})$
@@
$\lambda_{m}(\boldsymbol{\phi})=\beta^{2}\cdot|\xi_{m}(\boldsymbol{\phi})|^{2},\quad\text{for}\penalty 10000\ m=1:M$ (9)
@@
The Lagrangian function related to this doubly parameterized QCQP can be expanded into three terms:
@@
IV Solving the doubly variational QCQP/SDP
@@
Lemma 1 asserts that $F(\boldsymbol{\theta},\boldsymbol{\phi})$ is an expectation applied to the composite state of the two PQCs. This is important as it bypasses the need to measure each one of the numerous constraint expectations $\{F_{m}(\boldsymbol{\theta})\}_{m=1}^{M}$ of the OPF separately. This powerful feature was first identified in [51] for solving doubly variational SDPs on a quantum computer. In this work, we adopt this idea to the setting of solving QCQPs and SDPs over sparse graphs. We also propose a method for efficiently measuring the observables associated with sparse graphs. We uniquely leverage the underlying graph structure to efficiently measure the three observables using the two PQCs. We defer this discussion to Section V. If the observables can be measured efficiently, their gradients can be measured efficiently as well via the so-called parameter-shift rule (PSR). We review this rule and adapt it to the PD iterations in (14).
Let us focus on the gradient $\nabla_{\boldsymbol{\theta}}F_{0}(\boldsymbol{\theta}^{t})$. Suppose that the unitary matrix modeling PQCp takes the form
@@
The gradient of the expectation value in (16) with respect to $\boldsymbol{\phi}$ can be computed similarly by running PQCd by shifting the values of $\boldsymbol{\phi}^{t}$ for $2Q$ times or $2QS$ measurement samples. The gradient $\nabla_{\boldsymbol{\theta}}F(\boldsymbol{\theta}^{t},\boldsymbol{\phi}^{t})$ can be computed by running both PQCs while only shifting the entries of $\boldsymbol{\theta}^{t}$; and vice versa for computing the gradient $\nabla_{\boldsymbol{\phi}}F(\boldsymbol{\theta}^{t},\boldsymbol{\phi}^{t})$. It should be clear by now that, due to the PSR, evaluating gradients of expectations is equivalent to measuring expectations. Section V discusses how to measure the observables induced from a sparse graph.
Although the plain PD method has been extensively studied for convex/concave and nonconvex/concave saddle-point problems [31, 61, 35, 40], its convergence is not fully understood in the nonconvex/nonconcave setting, like the one in (13). Recently, a variation of the PD method, termed the extragradient (EG) method [34], has been shown to converge under certain conditions for nonconvex/nonconcave saddle-point problems [16]. Compared to the plain PD method, each EG iteration doubles the number of gradient evaluations of the Lagrangian function. More specifically, let vector $\mathbf{z}^{t}\coloneqq[\boldsymbol{\theta}^{t};\alpha^{t};\boldsymbol{\phi}^{t};\beta^{t}]$ collect the primal and dual variables at iterate $t$, and define the vector:
@@
for a step size $\mu_{z}^{t}>0$. In essence, the EG method first computes an intermediate point $\bar{\mathbf{z}}^{t}$ based on the plain PD method, and then updates $\mathbf{z}^{t}$ to $\mathbf{z}^{t+1}$ upon evaluating the gradient operator at the intermediate point. Despite doubling the number of gradient evaluations, EG iterates feature favorable convergence properties as established analytically in Section VI and corroborated numerically in Section VIII.
Before concluding this section, it is worth noting that, as with other VQAs, training the primal-dual PQC pair could be impeded by the so-called barren plateaus issue, according to which gradients vanish exponentially fast in terms of the number of qubits and depth of PQCs [45, 37]. This phenomenon can occur when designing VQAs for large-scale problems and/or using overly expressive PQCs. A related concern is that the PQC architectures that are both sufficiently expressive and free from barren plateaus are often efficiently simulated by classical algorithms [10]. Interestingly, recent results show that the so-called dynamic PQCs [15], which incorporate early measurements followed by measurement-conditional unitaries, and quantum recurrent embedding neural networks [30] can provably avoid barren plateaus while retaining high expressivity. Moreover, dynamic PQCs inherit the worst-case classical hardness of the universal unitary circuits, precluding the possibility of classical simulatability. Investigating the performance of the doubly variational QCQP/SDP on dynamic PQCs or quantum recurrent embedding neural networks for large-scale problems is an interesting future direction.
V Measuring Observables Induced by Sparse Graphs
This section reviews the extended Bell measurement (XBM) method proposed in [33] and explains how it can be adapted to efficiently measure observables induced from sparse graphs.
V.1 Extended Bell measurement (XBM) method
@@
$\braket{i|\mathbf{M}_{m}^{c}|j}=\begin{cases}\braket{i|\mathbf{M}_{m}|j},&\text{if}\penalty 10000\ i\oplus j=c\\ 0,&\text{otherwise}.\end{cases}$
@@
We next describe how to effect a measurement matrix $\mathbf{M}$ that uses a few colors.
V.2 Node permutation for qubit-efficient measurements